OpenStudy (anonymous):

whats the width of a rectangle that has a length of 6 cm and a perimeter of 16 cm.

OpenStudy (anonymous):

P = 2(length) + 2(width) 16 cm = 2 (6 cm) + 2 (width) 16 cm = 12 cm + 2 (width) 16cm - 12cm = 2 (width) 4cm = 2 (width) 4cm/2 = 2(width)/2 width = 2cm